. Skew rays are accepted into a large core diameter compared to the wavelength of thetransmitted light) step index fiber in air at a maximum axial angle of 450. Within the fiberthey change direction by 900at each reflection. Determine the acceptance angle formeridional rays or the fiber in air
1
Expert's answer
2020-11-11T09:00:08-0500

Given,

Axial angle "(\theta_{as}) =45^\circ"

As per the question, skew changes the direction by "90^\circ" at each reflection,

hence, "\gamma=45^\circ"

"\theta_{as}=\sin^{-1}(\frac{NA}{\cos\gamma})"

"\Rightarrow NA= \sin(\theta_{as})\times \cos(\gamma)"

"\Rightarrow NA = \sin(45)\times \cos(45^\circ)"

"\Rightarrow NA =\frac{1}{2}"

Now, Acceptance angle for meridional rays "\theta_{a}=\sin^{-1}(NA)"

"\Rightarrow \theta_{a}=\sin^{-1}(0.5)"

"\Rightarrow \theta_{a}=30^\circ"